# Out-of-equlibrium inference of stochastic model parameters through population data from generic consumer-resource dynamics

### v1

@@ III Model identifiability
$\mathcal{L}(\bm{\Theta}|M(\{t_{i}\}))=\prod_{i=1}^{T}P_{n_{A,i}}(\bm{\Theta},t_{i})$ (75)
@@
stands for an empirical data vector with $m$ observational abundances $n_{A,i}$ of the number of free predators. If the experiment only reports abundances at stationarity, these formulae apply to estimate parameters in the limit $t\to\infty$, given that $p_{A}(t)$ tends to the steady-state probability. Again, only the ratio $\nu/\alpha$ can be determined from abundance data at equilibrium. Model identifiability is illustrated with the plots of log-likelihood out-of and at equilibrium in Fig. 6, where only searching predators abundance data is used in ML parameter estimation. As in the first case, in which the abundances of each behavioral type were available, unique parameter estimates can be inferred from out-of-equilibrium model observations at different sampling times.
Figure 5: Negative log-likelihood profiles for the same synthetic matrix $M$ used for the heatmap in Fig. 3b. Negative log-likelihood profiles $\ell_{1}(\alpha)$ and $\ell_{2}(\nu)$ (left and right panels, respectively) yield confidence intervals as the abscissae of the intercept of the profiles $\ell_{1}(\alpha)$ and $\ell_{2}(\nu)$ with the horizontal dotted lines whose ordinates are equal to $\ell_{1}(\alpha^{\star})+\chi^{2}_{1,0.95}/2$ and $\ell_{2}(\nu^{\star})+\chi^{2}_{1,0.95}/2$, respectively. $(\alpha^{\star},\nu^{\star})$ refers to the point (marked with vertical, dotted lines) at which the maximum of Eq. (72) is reached.
@@
Figure 6: Heatmaps of negative log-likelihood using only free predator abundances. Left: representation of $-\log\mathcal{L}$, cf. Eq. (75) in the limit $t\to\infty$, obtained for a single model realization when samples are taken at equilibrium, using only free predators abundance information. Parameters $\alpha$ and $\nu$ are not identifiable. Right: a single out-of-equilibrium model replica yields a heatmap for negative log-likelihood —Eq. (75). The arrow indicates the actual pair $(\alpha,\nu)=(5,10)$ used for synthetic data generation, and the white point marks the minimum of $-\log\mathcal{L}$. Parameter values were chosen as in Fig. 3.
@@
Parameter estimation is illustrated with another in silico experiment based only on free predators abundance vectors —cf. Eq. (76). Figure 7 shows the distributions obtained in this case, gathering $3000$ ML parameter estimates for varying $n_{A}^{0}$, the overall number of consumers, and $T$, the number of observations (columns) used to construct vector $M$. Distributions for each parameter (upper panels) are very similar to those reported in Fig. 4, which took information of the full vector of abundances. Only for small-size samples (i.e., for $n_{A}^{0}$ and $T$ small), distributions and density plots (lower panels) appear to be more spread. This effect can be seen more clearly in confidence intervals, which are calculated using the empirical distributions in Fig. 7: reported intervals (Table 2), based only on free predator abundances, are wider than those listed in Table 1, for $n_{A}^{0}=20$ and $T=20$. Once the ML estimation procedure contains more information, as $n_{A}^{0}$ and $T$ increase, the intervals estimated from both synthetic experiments are consistently comparable.
@@ IV Discussion
In this contribution we have studied model identifiability, using a multi-resource consumer-resource dynamical model, and we have showed that, when independent model observations are taken at different sampling times, the model is truly identifiable in the sense that model parameters can be uniquely estimated from observations. On the contrary, if population numbers are observed at the same sampling time (this applies to the steady-state), then the model is set-identifiable because only a combination of parameters can be inferred from data. Synthetic data were used to calibrate the model in two in silico experiments: one in which only free predator population numbers are considered, and another in which the whole matrix of searching and handling predator numbers is available. If this model was to be calibrated using actual, experimental data, and independent observations were taken at stationarity (or at equal sampling times), the problem of non-identifiability would persist: the log-likelihood function would produce an infinite set of maxima, because the critical points conditions would reduce to one single equation —see the derivation of Eq. (96) in Appendix B.
@@
For stochastic processes other than the multi-resource consumer-resource model analyzed here, we can state that, as long as the distributions of observational variables depend on parameter combinations whose intersection is a manifold that does not reduce to a unique point, then non-identifiability issues will arise. This is particularly important at equilibrium, because many models are calibrated using steady-state distributions (which are the ones commonly available, unlike the out-of-equilibrium distributions, which are difficult to derive analytically). If those steady-state distributions depend on parameter ratios, for example, then the issue of parameter inference will arise. Because many quantities at equilibrium depend on dimensionless parameter combinations, we anticipate that the problem of model identifiability can be very frequent in stochastic models commonly used.
@@
Additional information should be available to conduct precise parameter inference at stationarity. When we take logarithms of Eq. (68), and change the sign, we obtain a log-likelihood function, which exactly corresponds to the one represented in the heat maps of Figure 3 as the two parameters vary between a minimum and a maximum value along the axes. As shown in Section II, the final expression for the joint stationary distribution (as well as its marginals) depends on the compound parameter controlled by the dimensionless ratio $\nu/\alpha$ (the relaxation rate in units of the attack rate). This non-dimensional quotient is the one that can be statistically inferred from data at equilibrium. If we were able to sort out, experimentally, individuals in the different behavioral modes at stationarity (attacking vs handling), we could identify, as shown in Figure 3, the true attack rate $\alpha$ on resources, respectively, provided we had external information to previously determine the relaxation rate $\nu$, and confidence intervals could be assigned (see Table 1). Regarding to the rest of parameters, this is, the total number of consumers, $n_{A}^{0}$, resource densitities $x_{R_{i}}$, and the volume, $N$, of the system, we believe they will be rather easy to be initially fixed and controlled in the setting of any experiment.
@@
Regarded to the generality of our results, the multinomial distribution derived in section II, either at stationarity or out of equilibrium, can be extended to a situation in which predator’s feeding efficiency varies non-linearly in response to both the same or different resources. Let $\alpha_{ij}$ ($i,j=1,\dots,S$) measure the effect of resource densities $x_{R_{i}}$ and $x_{R_{j}}$ on the formation of consumers handling on resource $i$. The effect of the density of the same resource or alternative resources on the formation of handling consumers of the focal resource $i$ can be taken into account, for example, if we assume the following $3\times S$ chemical reaction scheme:
@@
Proceeding in the same way as in [28], the functional response that arises from those fundamental processes corresponds to a multispecies version of Holling type III functional response. It is easy to see that, according to the reaction scheme (77)–(79) and because of mass action, the per capita consumption rate of a free/searching individual $X_{A}$ feeding on a focal resource $i$ becomes
@@
If only a single resource were present, the first term would correspond to a Holling type III functional response, with Hill exponent (i.e., the exponent to which resource densities are raised) $r=2$ [35]. The definition of per-capita consumption rates $\theta_{i}$ would be the only change in the derivation of the distributions, at stationarity or out of equilibrium, for this scheme of elemental processes. All the results derived for the multi-resource Holling type II case treated in section II hold in this more general case, because transition probability rates are exactly the same except for the replacement of parameter $\theta_{i}$ by the one in Eq. (80), which is still a constant parameter that depends on resource densities under the hypothesis of chemostatic conditions. Indeed, note that the two channels specified by Eqs. (77) and (78) can be subsumed into the transition probability rate for increasing the number of handling predators by one unit —first line of Eq. (6), with $\theta_{i}$ given by (80). Reaction (79) leads to the same transition rate for decreasing the number of handling predators by one unit —second line of Eq. (6). As a consequence, the previous derivation of the multi-resource Holling Type II stationary and out-of-equilibrium distributions applies here replacing parameter $\theta_{i}$ by the new definition. In both type II and III cases, the formulae that define probability parameters of the multinomial distributions remain valid with the replacement of $\theta_{i}$ by its expression in (80). Consequently, the experiments designed to carry out ML parameter estimation can be generalized to the Holling type III dynamics as well, together with the issues of parameter identifiability. Our results can be extended to more general consumer-resource contexts.
@@
We used the assumption of equal handling times in order to obtain closed-form solutions for the out-of-equilibrium joint distribution of population abundances. We expect the multinomial distribution to approximately hold when the distribution of handling times is not broad. In fact, we can always write $\nu_{j}=\nu+\epsilon e_{j}$ (with constant $\nu$), substitute this expression into Eq. (47) and neglect the term $\epsilon\sum_{j=1}^{S}\theta_{j}(1-z_{j})e_{j}$ if we assume that parameter $\epsilon$ is small enough compared to $\nu$. In principle, a perturbation series in this small parameter $\epsilon$ could be carried out in order to solve Eq. (48) for non-uniform rates $\nu_{j}$. Indeed, the closed-form solution we obtained in Section II for the PDE is the zero-th order term in the perturbation series expansion on $\epsilon$ of the general solution for non-equal relaxation rates. Future analytical work is needed to provide meaningful expressions for out-of-equilibrium distributions also when the differences in handling times are significant. On a side note, it is important to observe that the assumption of equal handling times does not impose any restriction in the case of one single resource: for the standard Holling type II consumer-resource feeding dynamics, the calculations that follow are fully general for the case $S=1$, and the multinomial out-of-equilibrium solution is exact in that case —provided that the dynamics starts with $n_{A}^{0}$ searching consumers initially.
@@
Classical feeding experiments have been performed for a variety of systems and taxa, where feeding rates have been measured mainly using the averaged expressions at fixed resource levels [21, 36, 37]. Here, we propose to use the distribution of behavioral types, thus allowing for refined estimates of parameters associated to elemental process. For the multi-resource Holling type II stochastic dynamics (as well as the extensions discussed here), we managed to give a complete analysis of the temporal dynamics, as well as its steady state. We described the potential inference methods to use to estimate, when identifiable, population parameters or characterized the subsets of the parameter space in which the model is set-identifiable. For models that are not fully solvable like the one studied here, one can always resort to numerical integration of the master equation in order to compute numerically the likelihood function and infer model parameters. We envision the application of the methods presented in this contribution in the design and realization of novel tests of predator-prey theory for controlled laboratory experiments [14] or field experiments [15, 11].
@@ Appendix A Derivation of the PDE satisfied by the generating function
$A:=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}$ $z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\left[T(\bm{n}|\bm{n}+\bm{e_{i}})P_{\bm{n}+\bm{e_{i}}}\right.$ (82)
$\left.-T(\bm{n}-\bm{e_{i}}|\bm{n})P_{\bm{n}}\right]$
@@
$B:=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}$ $z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\left[T(\bm{n}|\bm{n}-\bm{e_{i}})P_{\bm{n}-\bm{e_{i}}}\right.$ (83)
$\left.-T(\bm{n}+\bm{e_{i}}|\bm{n})P_{\bm{n}}\right],$
@@
$A$ $=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}[\nu_{i}(n_{i}+1)P_{\bm{n}+\bm{e}_{i}}-\nu_{i}n_{i}P_{\bm{n}}]$ (86)
×$\times z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}=\sum_{i=1}^{S}\nu_{i}(1-z_{i})\frac{\partial G}{\partial z_{i}}.$
@@ Appendix B ML estimation with equal sampling times
In this section we show that sampling data at equal times, $t_{1}=\dots=t_{T}=t$ ($t>0$), makes the model non-identifiable. In particular, this applies at equilibrium, where all times are sufficiently large (and therefore can be taken as equal). Here we write the multinomial distribution as
@@
Here we have defined the total resource density $\Lambda_{R}:=\sum_{i=1}^{S}x_{R_{i}}$. Then we can write $p_{i}=x_{R_{i}}h(\alpha,\tau)$, where
$h(\alpha,\tau)=\alpha\tau(1-e^{-t/\tau})$ (93)
and $\tau=(\nu+\alpha\Lambda_{R})^{-1}$ is the characteristic time.
@@
$\log\mathcal{L}(\alpha,\nu|M(t))=T\log C(\bm{n})+\sum_{k=1}^{T}\left\{n_{AR,k}\log h(\alpha,\tau)+n_{A,k}\log(1-\Lambda_{R}h(\alpha,\tau))+\sum_{i=1}^{S}n_{i,k}\log x_{R_{i}}\right\},$ (94)
where we have used that $p_{i}=x_{R_{i}}h(\alpha,\tau)$ and defined the $k$-th sampled total handing predators abundance as the sum $n_{AR,k}=\sum_{i=1}^{S}n_{i,k}$, and the $k$-th sampled number of free predators as $n_{A,k}=n_{A}^{0}-\sum_{i=1}^{S}n_{i,k}=n_{A}^{0}-n_{AR,k}$ —the last equality holds for every sample, $k=1,\dots,T$. Observe that the dependence on model parameters $\alpha$ and $\nu$ has been isolated in function $h$. Therefore, the first and last terms in the log-likelihood function do not contribute to derivatives with respect to those parameters.
@@
from which we can eliminate the function $h(\alpha,\tau)$ to obtain the condition to be satisfied by model parameters at the critical points. This defines a manifold in the parameter space, given by
@@
where we have defined $\langle n_{AR}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{AR,k}$ as the average across samples of handling consumers’ total abundance. The quantity $\langle n_{AR}\rangle_{T}$, relative to the product $\Lambda_{R}n_{A}^{0}$, determines only one single restriction to be satisfied by model parameters, in the $(\alpha,\nu)$ space. This immediately implies that the two parameters cannot be uniquely determined, yielding an infinite number of critical points satisfying Eq. (97), as we wanted to prove. In the case of exactly equal sampling times (either out of equilibrium or at stationarity), the model remains set-identifiable.
@@
as confirmed by the numerical optimization of the equilibrium log-likelihood function —here $\langle n_{A}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{A,k}=n_{A}^{0}-\langle n_{AR}\rangle_{T}$. The slope of the line of infinite maxima is determined by the average number of free predators, relative to the average total number of handling predators, times the overall density of resources. This is consistent with Eq. (70) if, at equilibrium, $\langle n_{A}\rangle$ and $\langle n_{AR}\rangle$ are estimated from data by averages across samples, i.e., by $\langle n_{A}\rangle_{T}$ and $\langle n_{AR}\rangle_{T}$, respectively.

### v2

@@ 4 Effect of steady-state and out-of-equilibrium sampling
$\mathcal{L}(\bm{\theta}|M(\{t_{j}\}))=\prod_{j=1}^{T}P_{n_{A,j}}(t_{j}|\bm{\theta})$ (78)
@@
stands for an empirical data vector with $T$ observational abundances $n_{A,j}$ of the number of free predators. We use the time-dependent log-likelihood to estimate $\alpha$ and $\nu$ depending on the sampling scheme used to generate the vector $M(\{t_{j}\})$. The ratio $\nu/\alpha$ can be inferred if the experiment only reports steady-state abundance data, but not rates themselves. Model identifiability is illustrated with the plots of log-likelihood out-of and at equilibrium in Fig. 5, where only searching predators abundance data is used for ML parameter estimation. As in the first experiment, in which the abundances of each behavioral type were available, unique parameter estimates can be inferred from out-of-equilibrium model observations.
@@
Figure 5: Heatmaps of negative log-likelihood using only free predator abundances. Representation of $-\log\mathcal{L}$, cf. Eq. (78) (i.e., using only free predators abundance information), in the $(\alpha,\nu)$ space, obtained for the same two sets of model samples generated for Fig. 2: true parameter values are $(\alpha,\nu)=(5,10)$ (arrow in the right panel), and stationary (out-of-equilibrium) samples yield the log-likelihood represented in the left (right) panel. Observe that the heatmaps only differ from those in Fig. 2 in the colorbar values, because here we are using only partial information to evaluate $-\log\mathcal{L}$ for the same model observations as in that figure. Again, we find parameter redundancy for stationary sampling, whereas a single minimum is obtained for out-of-equilibrium sampling (the white point marks the minimum of $-\log\mathcal{L}$). Remaining parameter values were chosen as in Fig. 1.
@@
Parameter estimation is illustrated with a second in silico experiment based only on free predators abundance vectors —cf. Eq. (79). Figure 6 shows the distributions obtained in this case, gathering $3000$ ML parameter estimates for varying $n_{A}^{0}$, the overall number of consumers, and $T$, the number of observations (columns) used to construct vector $M(\{t_{j}\})$. Distributions for each parameter (upper panels) are very similar to those reported in Fig. 3, which took information of the full vector of abundances. Only for small-size samples (i.e., for $n_{A}^{0}$ and $T$ small), distributions and density plots (lower panels) appear to be more spread in this case. This effect can be seen more clearly in confidence intervals, which are calculated using the empirical distributions in Fig. 6: reported intervals (Table 2), based only on free predator abundances, are wider than those listed in Table 1, for $n_{A}^{0}=20$ and $T=20$. Once the ML estimation procedure contains more information, as $n_{A}^{0}$ and $T$ increase, the intervals estimated from both synthetic experiments are consistently comparable.
@@ 5 Discussion
In this contribution we have studied model identifiability in two stochastic population models, and we have shown that, when independent model observations are taken at different sampling times in the transient regime, the model is truly identifiable in the sense that model parameters can be uniquely estimated from observations. On the contrary, if population numbers are observed at stationarity, then the model is set-identifiable because only a combination of parameters can be inferred from data. We derived model redundancy conditions using the generation function formalism in both a single-popultion birth-death-immigration model, and a multi-resource consumer-resource stochastic population model. For the latter, synthetic data were used to calibrate the model in two in silico experiments: one in which only free predator population numbers are considered, and another in which the whole matrix of searching and handling predator numbers is available. If this model was to be calibrated using actual, experimental data, and independent observations were taken at stationarity (or at equal sampling times), the problem of parameter redundancy would persist: the log-likelihood function would produce an infinite set of maxima, because the critical points conditions would reduce to one single equation —see the derivation of Eq. (99) in B.
@@
For stochastic processes other than the multi-resource consumer-resource model analyzed here, we can state that, as long as the distributions of observational variables depend on parameter combinations whose intersection is a manifold that does not reduce to a unique point, then identification failure will emerge. This is particularly important at equilibrium, because many models are calibrated using steady-state distributions (which are the ones commonly available, unlike the out-of-equilibrium distributions, which are difficult to derive analytically). If those steady-state distributions depend on parameter ratios, as in the examples discussed here, then the issue of parameter inference will arise. Because many quantities at equilibrium depend on dimensionless parameter combinations, we anticipate that the problem of parameter redundancy can be very frequent in stochastic models commonly used. This has important implications on ML parameter estimation using empirical data, because if the system under study is sampled once the steady-state has been reached, the problem of parameter redundancy is likely to arise.
@@
Additional information should be available to conduct precise parameter inference at stationarity. When we take logarithms of Eq. (74), and change the sign, we obtain a log-likelihood function, which exactly corresponds to the one represented in the heatmaps of Figure 2 as the two parameters vary between a minimum and a maximum value along the axes. As shown in Section 3, the final expression for the joint stationary distribution (as well as its marginals) depends on the compound parameter controlled by the dimensionless ratio $\nu/\alpha$ (the relaxation rate in units of the attack rate). This non-dimensional quotient is the one that can be statistically inferred from data at equilibrium. If we were able to sort out, experimentally, individuals in the different behavioral modes at stationarity (attacking vs handling), we could identify, as shown in Figure 2, the true attack rate $\alpha$ on resources, respectively, provided we had external information to previously determine the relaxation rate $\nu$, and confidence intervals could be assigned (see Table 1). Regarding to the rest of parameters, this is, the total number of consumers, $n_{A}^{0}$, resource densitities $x_{R_{i}}$, and the volume, $N$, of the system, we believe they will be rather easy to be initially fixed and controlled in the setting of any experiment.
@@
As for the generality of our results, the multinomial distribution derived in section 3, either at stationarity or out of equilibrium, can be extended to a situation in which predator’s feeding efficiency varies non-linearly in response to both the same or different resources. Let $\alpha_{ij}$ ($i,j=1,\dots,S$) measure the effect of resource densities $x_{R_{i}}$ and $x_{R_{j}}$ on the formation of consumers handling on resource $i$. The effect of the density of the same resource or alternative resources on the formation of handling consumers of the focal resource $i$ can be taken into account, for example, if we assume the following $3\times S$ chemical reaction scheme:
@@
Proceeding in the same way as in Palamara et al. (2022), the functional response that arises from those fundamental processes corresponds to a multispecies version of Holling type III functional response. It is easy to see that, according to the reaction scheme (80)–(82) and because of mass action, the per capita consumption rate of a free/searching individual $X_{A}$ feeding on a focal resource $i$ becomes
@@
If only a single resource were present, the first term would correspond to a Holling type III functional response, with Hill exponent (i.e., the exponent to which resource densities are raised) $r=2$ (Real, 1977). The definition of per-capita consumption rates $\theta_{i}$ would be the only change in the derivation of the distributions, at stationarity or out of equilibrium, for this scheme of elemental processes. All the results derived for the multi-resource Holling type II case treated in section 3 hold in this more general case, because transition probability rates are exactly the same except for the replacement of parameter $\theta_{i}$ by the one in Eq. (83), which is still a constant parameter that depends on resource densities under the hypothesis of chemostatic conditions. Indeed, note that the two channels specified by Eqs. (80) and (81) can be subsumed into the transition probability rate for increasing the number of handling predators by one unit —first equation of  (13), with $\theta_{i}$ given by (83). Reaction (82) leads to the same transition rate for decreasing the number of handling predators by one unit —second equation of (13). As a consequence, the previous derivation of the multi-resource Holling Type II stationary and out-of-equilibrium distributions applies here replacing parameter $\theta_{i}$ by the new definition. In both type II and III cases, the formulae that define probability parameters of the multinomial distributions remain valid with the replacement of $\theta_{i}$ by its expression in (83). Consequently, the experiments designed to carry out ML parameter estimation can be generalized to the Holling type III dynamics as well, together with the issues of parameter identifiability. Our results can be extended to more general consumer-resource contexts, unveiling the issue of parameter redundancy to be very common in stochastic population models.
@@
We used the assumption of equal handling times in order to obtain closed-form solutions for the out-of-equilibrium joint distribution of population abundances. By continuity, we expect the multinomial distribution to approximately hold when the distribution of handling times is not broad. In fact, we can always write $\nu_{j}=\nu+\epsilon e_{j}$ (for constant $\nu$), substitute this expression into Eq. (53) and neglect the term $\epsilon\sum_{j=1}^{S}\theta_{j}(1-z_{j})e_{j}$ if we assume that parameter $\epsilon$ is small enough compared to $\nu$. In principle, a perturbation series in this small parameter $\epsilon$ could be carried out in order to solve Eq. (54) for non-uniform rates $\nu_{j}$. Indeed, the closed-form solution we obtained in Section 3 for the PDE is the zero-th order term in the perturbation series expansion on $\epsilon$ of the general solution for non-equal relaxation rates. Future analytical work is needed to provide meaningful expressions for out-of-equilibrium distributions also when the differences in handling times are significant. On a side note, it is important to observe that the assumption of equal handling times does not impose any restriction in the case of one single resource: for the standard Holling type II consumer-resource feeding dynamics, the calculations that follow are fully general for the case $S=1$, and the multinomial out-of-equilibrium solution is exact in that case —provided that the dynamics starts with $n_{A}^{0}$ searching consumers initially.
@@
Classical feeding experiments have been performed for a variety of systems and taxa, where feeding rates have been measured mainly using the averaged expressions at fixed resource levels (Novak and Stouffer, 2021; Stouffer and Novak, 2021; DeLong, 2021). Here, we propose to use the distribution of behavioral types, thus allowing for refined estimates of parameters associated to elemental process. For the multi-resource Holling type II stochastic dynamics (as well as the extensions discussed here), we managed to give a complete analysis of the temporal dynamics, as well as its steady state. We described the potential inference methods to use to estimate, when identifiable, population parameters or characterized the subsets of the parameter space in which the model is set-identifiable. For models that are not fully solvable like those studied here, one can always resort to numerical integration of the master equation in order to compute numerically the likelihood function and infer model parameters. We envision the application of the methods presented in this contribution in the design and realization of novel tests of predator-prey theory for controlled laboratory experiments (Rosenbaum and Rall, 2018) or field experiments (Barbier et al., 2021; Abrams, 2022).
@@ Appendix A Derivation of the PDE satisfied by the generating function
$A:=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\left[T(\bm{n}|\bm{n}+\bm{e_{i}})P_{\bm{n}+\bm{e_{i}}}-T(\bm{n}-\bm{e_{i}}|\bm{n})P_{\bm{n}}\right]$ (85)
@@
$B:=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}\left[T(\bm{n}|\bm{n}-\bm{e_{i}})P_{\bm{n}-\bm{e_{i}}}-T(\bm{n}+\bm{e_{i}}|\bm{n})P_{\bm{n}}\right],$ (86)
@@
$A=\sum_{n_{1},\dots,n_{S}\geq 0}\sum_{i=1}^{S}[\nu_{i}(n_{i}+1)P_{\bm{n}+\bm{e}_{i}}-\nu_{i}n_{i}P_{\bm{n}}]z_{1}^{n_{1}}\cdots z_{S}^{n_{S}}=\sum_{i=1}^{S}\nu_{i}(1-z_{i})\frac{\partial G}{\partial z_{i}}.$ (89)
@@ Appendix B ML estimation with equal sampling times
In this section we show that sampling data at equal times, $t_{1}=\dots=t_{T}=t$ ($t>0$), makes the model unidentifiable. Here we write the multinomial distribution as
@@
Here we have defined the total resource density $\Lambda_{R}:=\sum_{i=1}^{S}x_{R_{i}}$. Then we can write $p_{i}=x_{R_{i}}h(t;\alpha,\tau)$, where
$h(t;\alpha,\tau)=\alpha\tau(1-e^{-t/\tau})$ (96)
and $\tau=(\nu+\alpha\Lambda_{R})^{-1}$ is the characteristic time scale.
@@
$\log\mathcal{L}(\alpha,\nu|M(t))=T\log C(\bm{n})+\sum_{k=1}^{T}\left\{n_{AR,k}\log h(t;\alpha,\tau)+n_{A,k}\log(1-\Lambda_{R}h(t;\alpha,\tau))+\sum_{i=1}^{S}n_{i,k}\log x_{R_{i}}\right\},$ (97)
where we have used that $p_{i}=x_{R_{i}}h(t;\alpha,\tau)$ and defined the $k$-th sampled total handing predators abundance as the sum $n_{AR,k}=\sum_{i=1}^{S}n_{i,k}$, and the $k$-th sampled number of free predators as $n_{A,k}=n_{A}^{0}-\sum_{i=1}^{S}n_{i,k}=n_{A}^{0}-n_{AR,k}$ —the last equality holds for every sample, $k=1,\dots,T$. Observe that the dependence on model parameters $\alpha$ and $\nu$ has been isolated in function $h$. Therefore, the first and last terms in the log-likelihood function do not contribute to derivatives with respect to those parameters.
@@
from which we can eliminate the function $h(t;\alpha,\tau)$ to obtain the condition to be satisfied by model parameters at the critical points. This defines a manifold in the parameter space, given by
@@
where we have defined $\langle n_{AR}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{AR,k}$ as the average across samples of handling consumers’ total abundance. The quantity $\langle n_{AR}\rangle_{T}$, relative to the product $\Lambda_{R}n_{A}^{0}$, determines only one single restriction to be satisfied by model parameters, in the $(\alpha,\nu)$ space. This immediately implies that the two parameters cannot be uniquely determined, yielding an infinite number of critical points satisfying Eq. (100), as we wanted to prove. In the case of exactly equal sampling times (either out of equilibrium or at stationarity), the model remains set-identifiable. At least two sampling times must be distinct in order for $\alpha$ and $\nu$ being non-redundant.
@@
as confirmed by the numerical optimization of the equilibrium log-likelihood function —here $\langle n_{A}\rangle_{T}:=\frac{1}{T}\sum_{k=1}^{T}n_{A,k}=n_{A}^{0}-\langle n_{AR}\rangle_{T}$. The slope of the line of infinite maxima is determined by the average number of free predators across samples, relative to the average total number of handling predators, times the overall density of resources. This is consistent with Eq. (76) if, at equilibrium, $\langle n_{A}\rangle$ and $\langle n_{AR}\rangle$ are estimated from data by their corresponding averages across samples, i.e., by $\langle n_{A}\rangle_{T}$ and $\langle n_{AR}\rangle_{T}$, respectively.